OpenStudy (nightmarenight):

Will fan and give medal???? Which set of side lengths is from a right triangle? 4, 8, 10 9, 10, 15 6, 9, 14 21, 28, 35

OpenStudy (bibby):

\(a^2+b^2 > c^2\) acute \(a^2+b^2 < c^2\) obtuse \(a^2+b^2 = c^2\) right

OpenStudy (bibby):

where c is the longest side and a, b are the legs

OpenStudy (nightmarenight):

wat @bibby

OpenStudy (bibby):

for example, let's check it 4,8,10 is right \(4^2+8^2=16+64=80\) 10 is the hypotenuse. \(10^2=100\) so as we can see \(a^2+b^2<c^2\) and so it isn't right

OpenStudy (nightmarenight):

So A isn't the answer

OpenStudy (bibby):

yeah. I have a good feeling about 21, 28, 35. test it out

OpenStudy (nightmarenight):

\[21^{2}+28^{2}=441+784=1225 35^{2}=\]

OpenStudy (nightmarenight):

that came out wrong, one sec

OpenStudy (nightmarenight):

21^2 + 28^2 = 441 + 784 = 1225 and 35^2 = 1225

OpenStudy (bibby):

so it would seem that \(a^2+b^2=c^2\). there she is

OpenStudy (nightmarenight):

So it's D c:

OpenStudy (bibby):

yeah, you can test t he other 2 to make sure if you'd like

OpenStudy (nightmarenight):

I have a lot to get done tonight, I trust that it's D xD thank you!

OpenStudy (bibby):

any time :)